OpenStudy (anonymous):

How would you solve for the lim x-> 0 of (integral 1 to 1+x of (cos t / t) dt)/x ?

OpenStudy (lgbasallote):

\[\lim_{x \rightarrow 0} \int_{1}^{1+x} \frac{\cos t}{t} dt\] right??

OpenStudy (anonymous):

The integral is the numerator with x as a denominator*

OpenStudy (lgbasallote):

uhh complicated..i'll call the os genius @satellite73 :)

OpenStudy (lgbasallote):

he'll LOVE that hehe

OpenStudy (anonymous):

i think you can use l'hopital

OpenStudy (anonymous):

I would do that, but I don't know how to get the derivative of the integral in this case

OpenStudy (anonymous):

\[f(x)=\int_1^{1+x}\frac{\cos(t)}{t}dt\] \[g(x)=x\] and you want \[\lim_{x\to 0}\frac{f(x)}{g(x)}\]

OpenStudy (anonymous):

derivative of the integral is the integrand so you get \[f'(x)=\frac{\cos(1+x)}{x}\], \[g'(x)=1\]

OpenStudy (anonymous):

now take the limit as x goes to zero, and hmmm something is wrong

OpenStudy (anonymous):

oh my derivative is wrong!

OpenStudy (anonymous):

I know the answer is cos 1, if that helps. I just don't understand why/how.

OpenStudy (anonymous):

\[f'(x)=\frac{\cos(1+x)}{1+x}\]

OpenStudy (anonymous):

i forgot to replace the \(t\) in the integrand by \(1+x\) in the denominator as well as in the numerator now take \[\lim_{x\to 0}\frac{\cos(1+x)}{1+x}=\cos(1)\]

OpenStudy (anonymous):

does the lower bound of the integral become irrelevant in this type of problem?

OpenStudy (anonymous):

only part that requires thinking is taking the derivative of \[f(x)=\int_1^{1+x}\frac{\cos(t)}{t}dt\] using the chain rule

OpenStudy (anonymous):

you mean for taking the derivative? yes, the lower bound is irrelevant if it is a number and does not include a function of \(x\)

OpenStudy (anonymous):

*if it does not include a function of \(x\)

OpenStudy (anonymous):

the lower bound determines the constant, but the derivative of a constant is 0

OpenStudy (anonymous):

makes sense, thank you!

OpenStudy (anonymous):

yw